OpenStudy (emac):

Use implicit differentiation to find the slope of the tangent line to the curve xy^3+xy=10 at the point (5,1)

OpenStudy (emac):

I already did the implicit differentiation and got \[\frac{ -y^3+y }{ 3xy^2+x }\]

OpenStudy (irishboy123):

check the minus sign on the top of the fraction

OpenStudy (emac):

Should it be positive?

OpenStudy (irishboy123):

\[\dfrac{dy}{dx} = -\dfrac{ y^3+y }{ 3xy^2+x }\]

OpenStudy (emac):

Okay gotcha, I'm just not sure where to go from here

OpenStudy (irishboy123):

plug in x and y ??

OpenStudy (emac):

ohh okay i using my original equation which was giving me the wrong answer but I got it now! Thanks!!

OpenStudy (irishboy123):

try out the Implicit Function Theorem: \(\dfrac{dy}{dx} = -\dfrac{f_x}{f_y}\) :-))